If In 2=0.6931, In 5 1.6094, and In 9 = 2.1972, find each of the following: a. In 45 Preview b. In 18 Preview review 6) Preview 9 e. In 4 Preview f. In 81 review g. In 125 = revieW h. In 16 = Preview
Answer #1

a) ln 45 = ln( 5 x 9) = ln 5 + ln 9 =1.6094 + 2.1972 = 3.8066

b) ln 18 = ln(2 * 9) = ln 2 + ln 9 = 0.6931+ 2.1972 = 2.8903

c) ln(1/5) = ln 1 - ln 5 = 0 -ln 5 = -1.6094

d) ln(1/9) = ln 1 - ln 9 = 0 -ln 9 = -2.1972